When the sum of twice the number and three is subtracted from 6, the result is equal to ten minus that number. Find the number.

Answer by DrBeeee(684)

Let n = the number.
The problem statement reduces to the equation
(1) 6 - (2n+3) = 10 -n
Now solve for n
(2) 6 - 2n - 3 = 10 - n or
(3) 3 - 10 = 2n - n or
(4) -7 = n or
(5) n = -7 Check this using (1).
Is (6-(2*(-7)+3) = 10 -(-7))?
Is (6-(-14+3) = 10 + 7)?
Is (6-(-11) = 17)?
Is (6+11 = 17)?
Is (17 = 17)? Yes
Answer: the number is -7.